A circle representing a pool is graphed with a center at the origin. Grant enters the pool at point A and swims over to a friend who is located at point B.

Which equation represents Grant’s path?
y = 2 – 4x
y = 4 –
y = 6 –
y = 8 – 2x
1
Expert's answer
2020-02-18T07:21:00-0500

Point A: (8,0),

Point B: (-4,6)

Slope: "\frac{6-0}{-4-8}=-\frac{1}{2}"

Grant's path: "y=-\frac{1}{2}(x-8) \;\;or\;\;y=4-\frac{1}{2}x"